OpenStudy (anonymous):

Fiber-optic cables are used widely for internet wiring, data transmission, and surgeries. When light passes through a fiber-optic cable, its intensity decreases with the increase in the length of the cable. If 1500 lumens of light enters the cable, the intensity of light decreases by 3.4% per meter of the cable. Part A: Can this situation be represented by a linear function? Justify your answer. Part B: Write a function f(x) to represent the intensity of light, in lumens, when it has passed through x meters of the cable.

OpenStudy (anonymous):

Part C: Some scientists are trying to make a cable for which the intensity of light would decrease by 5 lumens per unit length of the cable. Can this situation be represented by a linear function? Justify your answer and write the appropriate function to represent this situation if 1500 lumens of light enter the cable.

OpenStudy (anonymous):

Can someone help me with these

OpenStudy (anonymous):

@zepdrix @dan815

OpenStudy (anonymous):

@tester97 @mathmale

OpenStudy (anonymous):

@jcpd910

OpenStudy (anonymous):

Okay, here we go. y = mx + b y = -51 + 1500 The slope is -51 because our lumens are decreasing from 1500, and because we started with 1500, our y intercept is 1500. And: f(m) = 1500 - 1500(0.034)m

OpenStudy (anonymous):

For part C, we are creating another linear equation. If our slope is -5, where it decreases 5 lumens per unit length, then our equation is: y = -5 + 1500

OpenStudy (anonymous):

Remember, it's written as y = -51x + 1500 and y = -5x + 1500

OpenStudy (anonymous):

because we need our x coordinate to have a definition.

OpenStudy (anonymous):

ok gotchu